Given that 𝑧1 = 3 + 𝑖 π‘Žπ‘›π‘‘ 𝑧2 = 2 βˆ’ 𝑖: i. Find the modulus and argument of 𝑧1/𝑧2 (5 marks) ii. Express 𝑧1/𝑧2 in polar and exponential form (2 marks) iii. Use de Moivre’s theorem to find an expression for ( 𝑧1/𝑧2) ^4Β 


1
Expert's answer
2021-06-07T17:20:31-0400

"i)\\
|\frac{z_1}{z_2}|=|\frac{3+i}{2-i}|=|\frac{(3+i)(2+i)}{2^2-i^2}|=|\frac{5(1+i)}{5}|=|1+i|=\sqrt{1^2+1^2}=\sqrt2\\
arg(\frac{z_1}{z_2})=arg(1+i)=tan^{-1}(1)=\frac{\pi}{4}\\
ii)\\
Since, z=re^{i\theta}.\\
\frac{z_1}{z_2}=1+i=\sqrt2e^{i\frac{\pi}{4}}.\\
\text{This is the exponential form.}\\
Polar form, z=re^{i\theta}=r(cos\theta +isin\theta).\\
\frac{z_1}{z_2}=\sqrt2e^{i\frac{\pi}{4}}=\sqrt2({cos\frac{\pi}{4}}+isin{\frac{\pi}{4}})\\
iii)\\
\text{by using de Moivre's theorem,}\\
(\frac{z_1}{z_2})^4=(\sqrt2)^4e^{i4\frac{\pi}{4}}
=4e^{i\pi}
=4(cos\pi+isin\pi)=4((-1)+i(0))=-4"